A 210-ton train has its speed reduced from 40 to 30 miles per hour in 0.50 minute. If the whole of the work done against frictional resistance of the brakes converted into heat, find the heat developed in BTU.


1
Expert's answer
2021-11-26T16:52:02-0500
"Q=0.5(210000)(17.88^2-13.41^2)\\Q=1.47\cdot10^7\ J=1.4\cdot10^4\ Btu"